OpenStudy (anonymous):

differentiate using quotient rule (x-1)(x+2)/x

OpenStudy (anonymous):

\[ \frac{(x-1)(x+2)}{x}=\frac{x^2+x-2}{x}\\ \left(\frac f g \right)'=\frac{gf'-fg'}{g^2}\\ f=x^2+x-2,\ g=x\\ f'=2x+1,\ g'=1 \]

OpenStudy (anonymous):

You can take it from there?

OpenStudy (anonymous):

yh it will be x(2x+1)-(x^2+x-2)1/x^2

OpenStudy (anonymous):

s this ryt nbouscal

OpenStudy (anonymous):

2x^2+x-x^2-x+2/x^2

OpenStudy (anonymous):

Can you simplify that further?

OpenStudy (anonymous):

yh will it be \[x^2+ 2/x^2\]

OpenStudy (anonymous):

Yeah, which you can simplify further to \(1+\dfrac{2}{x^2}\)

OpenStudy (anonymous):

and this = 2 am i ryt ?

OpenStudy (anonymous):

k s this the final answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thk u so mch